OpenStudy (anonymous):

log (6x + 2) = 8 solve for x

OpenStudy (anonymous):

10^log(6x+2)=10^8 6x+2=10^8 6x=10^8-2 x=(10^8-2)/6

OpenStudy (anonymous):

what base is the log ?

OpenStudy (anonymous):

base 10 ?

OpenStudy (anonymous):

I was assuming it was base 10

OpenStudy (anonymous):

If I had your dp, I'd assume it was base e :)

OpenStudy (aravindg):

ya

OpenStudy (anonymous):

wolfram assumes base e , , but since the question was log and not ln,

OpenStudy (aravindg):

giv medal plzzz

OpenStudy (anonymous):

wolfram converts ln --> log :) http://www.wolframalpha.com/input/?i=ln%28e%29

OpenStudy (anonymous):

yeah, it is of base ten